# SLOctolyzer: Fully automatic analysis toolkit for segmentation and feature extracting in scanning laser ophthalmoscopy images

### v1

@@ Abstract
Purpose To describe SLOctolyzer: an open-source analysis toolkit for en face retinal vessels appearing in infrared reflectance scanning laser ophthalmoscopy (SLO) images.
Methods SLOctolyzer includes two main modules: segmentation and measurement. The segmentation module use deep learning methods to delineate retinal anatomy, while the measurement module quantifies key retinal vascular features such as vessel complexity, density, tortuosity, and calibre. We evaluate the segmentation module using unseen data and measure its reproducibility.
Results SLOctolyzer’s segmentation module performed well against unseen internal test data (Dice for all-vessels, 0.9097; arteries, 0.8376; veins, 0.8525; optic disc, 0.9430; fovea, 0.8837). External validation against severe retinal pathology showed decreased performance (Dice for arteries, 0.7180; veins, 0.7470; optic disc, 0.9032). SLOctolyzer had good reproducibility (mean difference for fractal dimension, -0.0007; vessel density, -0.0003; vessel calibre, -0.3154 $\mu$m; tortuosity density, 0.0013). SLOctolyzer can process a macula-centred SLO image in under 20 seconds and a disc-centred SLO image in under 30 seconds using a standard laptop CPU.
@@
Translational Relevance SLO images are captured simultaneous to optical coherence tomography (OCT), and we believe our software will be useful for extracting retinal vascular measurements from large OCT image sets and linking them to ocular or systemic diseases.
@@ 1 Introduction
The retina is a highly vascularised light-sensitive tissue at the back of the eye. The retinal microvasculature branches from the optic nerve head along the inner surface of the retina and is commonly imaged using colour fundus photography, which typically covers a 35–50 degree field of view around the posterior pole.
@@
Confocal near infrared reflectance scanning laser ophthalmoscopy (SLO) utilise long wavelengths of around 820nm and also captures an en face view of the superficial retinal vessels 1. SLO images are typically captured simultaneous to optical coherence tomography (OCT) scans of the inner retinal layers (supplementary Fig. S1) acting as a localiser to position the OCT scan at the back of the eye accurately and typically images a restricted field of view of 30 degrees (approximately 9 mm2). Additionally, confocal imaging leverages laser-point raster scanning to prevent stray light interference during acquisition 2, thus providing higher contrast visualisation of the en face retinal vasculature over regular colour fundus images. Moreover, the interferometry of OCT imaging enables an exact conversion from pixel to physical space, a crucial step in forming clinically meaningful datasets of features of the en face retinal vessels 3. In colour fundus images, heuristic approaches are commonly used to obtain physical measurements, such as optic disc area normalisation 4. However, despite the detail SLO images capture they are often discarded due to the primary aim of OCT imaging to visualise the cross-sectional retinal layers.
@@
This improved visualisation, increased accuracy in measurement and growing abundance from OCT capture make the SLO imaging modality an exciting frontier of en face retinal vessel analysis. Manual delineation of en face retinal vessels is prohibitively time-consuming, labour-intensive, and prone to inaccuracies. Consequently, extensive research has focused on segmenting retinal vessels, primarily in the colour fundus modality 5, with less emphasis on SLO images.
@@
At the time of writing, very few methods have been reported for retinal vessel analysis for SLO images 6; 7; 8; 9; 10, and none for other major landmarks on the SLO image, the fovea and optic disc. Additionally, none of these methods have been made available to the research community, which has the potential to impact the standardisation of SLO-derived retinal vascular parameters.
@@
Equipping automatic imaging methods with the tools for measuring downstream features for research and clinical use is an important element of any research pipeline 11; 12; 13; 14; 15. This allows the general researcher to use these methods in an end-to-end manner, and analyse their own retinal images without specialist training or a background in image processing. Unfortunately, there is a lack of similar pipelines for SLO images.
@@
Accordingly, we aimed to develop and release a fully automatic analysis toolkit, SLOctolyzer, for segmentation and subsequent measurement of retinal vessels as seen on SLO images. Here, we introduce and quantitatively validate the segmentation models which detect the binary vessels, and classification of arteries and veins, as well as detection of the optic disc and fovea.
@@
  • SLOctolyzer supports manual annotation to correct erroneous segmentations.
@@ 2 Methods
Fig. 1 summarises the key elements of SLOctolyzer’s pipeline, which contains a segmentation module and a measurement module (of which the latter is based on features and code produced by Automorph 12). We have released SLOctolyzer in a frictionless manner so that it can be used without author permissions, proprietary software, or specialist training. SLOctolyzer is freely available here: https://github.com/jaburke166/SLOctolyzer.
@@ 2.1 Data
Data used to build the segmentation module include RAVIR 10, a publicly available dataset of pathological SLO images with the arteries and veins segmented, and three additional cohorts related to systemic health: i-Test 16, a cohort of pregnant women at late gestation who are experiencing either a normative, pre-eclamptic or small baby pregnancy; PREVENT-Dementia 17; 18, a cohort of mid-life individuals, half of whom are at genetic risk of developing later life dementia; FutureMS 19; 20, a cohort of individuals with newly diagnosed relapsing-remitting multiple sclerosis (MS). All studies and cohorts adhered to the Declaration of Helsinki and received relevant ethical approval and informed consent from all subjects was obtained in all cases from the host institution.
@@
Table 1 summarises image characteristics of each cohort. All SLO images were captured using a spectral domain Heidelberg SPECTRALIS HRA+OCT imaging device (Heidelberg Engineering, Heidelberg, Germany). All images were either centred on the macula or the optic disc. SLO images from Heidelberg HRA+OCT imaging devices are grayscale, capturing a 30 degrees field of view (approximately 9 mm2) using a laser light with wavelength 820nm. All cohorts used a HRA+OCT SPECTRALIS Standard Module except for the i-Test study, which used a portable version, the SPECTRALIS FLEX, for 24 participants out of the 93 available. Dependent upon settings prior to acquisition, SLO images from Heidelberg devices have either a pixel resolution of 768 $\times$ 768 or 1536 $\times$ 1536 (pixel height $\times$ width).
@@
Study Participants Images Right eyes Retinal pathology HRA+OCT Module Image resolution Location
RAVIR 10 23 23 14 Yes Standard 768 Disc
Healthy 21 7 7 7 No Standard 1536 Disc
PREVENT 17; 18 144 285 142 No Standard 1536 Disc
i-Test 16 93 186 93 No Standard & FLEX 768 Macula
FutureMS 19; 20 15 15 9 No Standard 1536 Disc
@@
Table 1: Image characteristics of the five cohorts used to build SLOcolyzer’s segmentation module. Image resolution is in pixels (for both lateral and axial directions), location refers to the centring of the scan, i.e. if it’s macula-centred disc-centred. HRA, Heidelberg retina angiography; OCT, optical coherence tomography.
@@
SLOctolyzer’s segmentation module contains three segmentation models, one for binary vessel detection, one for fovea detection and another for artery-vein and optic disc detection. Table 2 outlines which cohorts and how many eyes/images of the cohorts each model used for training and evaluation. For all studies except for FutureMS, we selected all available participants for modelling. For FutureMS, a subset of 15 participants from the wider cohort 20 were selected. These images were chosen based on image features relevant for vessel segmentation, such as blur, illumination, contrast, and abnormal features like vessel tortuosity or optic nerve head atrophy (supplementary Fig. S2). Similarly, for the i-Test cohort, 15 participants were selected from the wider cohort for building the artery-vein-optic disc (AVOD) model.
@@ 2.2.1 Binary vessel detection
A total of 30 SLO images (RAVIR & Healthy) were used to train and evaluate the segmentation model for binary vessel detection. 23 of these were sourced from the RAVIR dataset, of which manual pixel-level annotation of the arteries and veins were done by the same, experienced retinal image grader 10. Segmentation over the optic nerve head was only performed if artery and vein classes were able to be resolved. Vessel labels were defined as both the artery and vein class combined. The remaining seven SLO images were made available from healthy eyes and were manually segmented by an experienced image analyst (author S.G.) using ITK-Snap 22, which supports pixel-level annotation.
@@ 2.2.2 Fovea detection
Fovea detection was performed by the same experienced grader (author J.B.) and used all 516 SLO images made available.
For the i-Test cohort, access to both the SLO and OCT was possible, which allowed the fovea coordinate on the SLO to be cross-referenced using it’s location on the corresponding fovea-centred OCT B-scan. The fovea location on each OCT B-scan was detected using Choroidalyzer 15, a fully automatic toolkit for choroid analysis and fovea detection in OCT images. For the remaining images, only the SLO was available and the fovea coordinate was selected manually, using an in-house GUI written in Python (version 3.11.6).
We assessed intra-rater variability by one user (author J.B.) through re-selecting the fovea pixel coordinate for all 516 images, two months after the initial manual grading, and the results of which can be found in supplementary Table S1. Repeatability against the initial i-Test fovea coordinates acted as comparison to ground truth, and the remaining images were used to assess intra-rater variability. Intra-rater variability had an average intra-class correlation (ICC(3,1)) of 0.9934 for both $x$ and $y$ coordinates, and comparison to ground truth scored an average ICC of 0.8206 across both $x$ and $y$ coordinates.
@@ 2.2.3 Artery-vein-optic disc detection
Three experienced image analysts (authors J.B., S.G., A.T.) were given 14 SLO images to manually annotate the arteries, veins, and optic disc. For each SLO image, a binary vessel map was generated using the binary vessel detection model. Each grader then used ITK-Snap 22 for anatomical segmentation of arteries, veins and the optic disc. See supplementary Fig. S3 and S4 for the vessel labelling rules and segmentation protocol each rater followed for consistent and comparative segmentation.
Each set of 14 images had two images with distinct overlap between pairwise graders. Thus, there were a total of 30 unique SLO images used for modelling: 15 from the i-Test cohort and 15 from the FutureMS cohort. This overlap allowed for a total of six SLO images which could be used for comparison. We calculated inter-rater agreement for all-vessel, artery, vein, and optic disc segmentation using the Dice coefficient, whose results can be found in supplementary Table S2. All values were found to be greater than 0.93, suggesting an excellent degree of consistency between graders.
@@ 2.2.4 The RAVIR dataset
The RAVIR dataset 10 was intended to be used for training the artery-vein-optic disc (AVOD) model, as it contains suitable examples of severe retinal pathology. However, before modelling we reviewed the segmentation labels and observed inconsistent artery and vein classification, an example of which can be seen in supplementary Fig. S6. Therefore, we decided to use the RAVIR dataset as an external test set for the AVOD model, after correcting for any misclassified veins or arteries. This external test set presents a significant challenge to the AVOD model as it contains examples of severe retinal pathology.
@@ 2.3 Segmentation module
We used python 3.12, PyTorch 2.0, Segmentation Models PyTorch 23 and MONAI 24.
@@ 2.3.1 Binary vessel detection
The model and training pipeline has been discussed previously 25. Briefly, the model was a custom U-Net deep learning architecture trained from scratch on a total of 30 SLO images. The dataset was split randomly into training (n=24), validation (n=2) and test (n=4) sets. Offline data augmentation randomly cropped 20 patches of size 320 $\times$ 240 to artificially increase the training set, reduce computational load and accelerate the training rate.
The model was trained to minimise the Dice Focal loss 24, using the Adam 26 optimiser (with default optimiser hyperparameters) for 600 epochs using a batch size of 20. The initial learning rate was set as 1e-3 until epoch 300, at which point the learning rate was reduced to 1e-4 for the remainder of training.
@@ 2.3.2 Fovea detection
Fovea segmentation is a significantly easier task than vessel detection or artery-vein classification, so our model used a pre-trained 23, lightweight U-Net deep learning architecture with a MobileNetV3 27 backbone, previously trained using ImageNet 28. This model used all available images from all cohorts, comprising 516 SLO images.
The dataset was split into training ($n=359$), validation ($n=104$) and test sets ($n=54$) at the participant-level, so there was no overlap of participants between sets. All images were resized to a common pixel resolution of $768\times 768$ for training. To facilitate learning through segmentation, training labels were circular binary masks whose centre was the fovea coordinate with radius of 60 pixels.
The model was trained to minimise the binary cross entropy loss using the AdamW 29 optimiser for 200 epochs using a batch size of 8. During training, we applied two sets of random augmentations per image, per epoch, artificially doubling the batch size. An initial learning rate of 5e-4 was set, and after a 10-epoch linear warm up to a learning rate of 5e-3, it decayed following a cosine relationship for 90 epochs, with the learning rate set at 5e-4 for the remaining 100 epochs.

### v2

@@ Abstract
Purpose The purpose of this study was to introduce SLOctolyzer: an open-source analysis toolkit for en face retinal vessels in infrared reflectance scanning laser ophthalmoscopy (SLO) images.
Methods SLOctolyzer includes two main modules: segmentation and measurement. The segmentation module uses deep learning methods to delineate retinal anatomy, and detects the fovea and optic disc, whereas the measurement module quantifies the complexity, density, tortuosity, and calibre of the segmented retinal vessels. We evaluated the segmentation module using unseen data and measured its reproducibility.
Results SLOctolyzer’s segmentation module performed well against unseen internal test data (Dice for all-vessels = 0.91; arteries = 0.84; veins = 0.85; optic disc = 0.94; and fovea = 0.88). External validation against severe retinal pathology showed decreased performance (Dice for arteries = 0.72; veins = 0.75; and optic disc = 0.90). SLOctolyzer had good reproducibility (mean difference for fractal dimension = -0.001; density = -0.0003; calibre = -0.32 microns; and tortuosity density = 0.001). SLOctolyzer can process a 768 $\times$ 768 pixel macula-centred SLO image in under 20 seconds and a disc-centred SLO image in under 30 seconds using a laptop CPU.
@@
Translational Relevance SLO images are captured simultaneous to optical coherence tomography (OCT), and we believe SLOctolyzer will be useful for extracting retinal vascular measurements from large OCT image sets and linking them to ocular or systemic diseases.
@@ 1 Introduction
The retina is a highly vascularised light-sensitive tissue at the back of the eye. The retinal microvasculature branches from the optic nerve head along the inner surface of the retina and is commonly imaged using colour fundus photography, which typically covers a 35- to 50-degree field of view around the posterior pole.
@@
Confocal near infrared reflectance scanning laser ophthalmoscopy (SLO) utilise long wavelengths of around 820nm and also captures an en face view of the superficial retinal vessels 1. SLO images are typically captured simultaneous to optical coherence tomography (OCT) scans of the inner retinal layers (supplementary Fig. S1) acting as a localiser to position the OCT beam at the back of the eye accurately and typically images a restricted field of view of 30-degrees. Additionally, confocal imaging leverages laser-point raster scanning to prevent stray light interference during acquisition 2, thus providing higher contrast visualisation of the en face retinal vasculature over regular colour fundus images.
The interferometry of OCT imaging allows for micron-level axial depth visualisation on the cross-section which is independent of biometric factors like axial length and corneal or refractive properties of the eye 3. When these parameters are accounted for, the transverse plane can also be imaged with micron-level precision, enabling standardised physical measurements of the en face retinal vessels using microns-per-pixel conversion factors 4. In contrast, colour fundus imaging typically relies on heuristic methods, such as optic disc area normalisation, to achieve standardised measurements 5. This suggests that SLO imaging could potentially provide more accurate and clinically relevant measurements of en face retinal vessels compared to traditional fundus imaging, given known biometric factors of the eye are accounted for. Nevertheless, despite the high level of detail captured by SLO images, they are often overlooked, as the primary focus of OCT imaging remains on visualising the cross-sectional retinal layers.
@@
This improved visualisation, possibly increased accuracy in measurement and growing abundance from OCT capture make the SLO imaging modality an exciting frontier of en face retinal vessel analysis. Manual delineation of en face retinal vessels is prohibitively time-consuming, labour-intensive, and prone to inaccuracies. Consequently, extensive research has focused on segmenting retinal vessels, primarily in the colour fundus modality 6, with less emphasis on SLO images.
@@
At the time of writing, very few methods have been reported methods for retinal vessel analysis on SLO images 7; 8; 9; 10; 11, and none for other major landmarks on the SLO image such as the fovea and optic disc. Additionally, none of these methods have been made available to the research community, which has the potential to impact the standardisation of SLO-derived retinal vascular metrics.
@@
Equipping automatic imaging methods with the tools for measuring downstream features for research and clinical use is an important element of any research pipeline 12; 13; 14; 15; 16. This allows the general researcher to use these methods in an end-to-end manner, and analyse their own retinal images without specialist training or background in image processing. Unfortunately, there is a lack of similar pipelines for SLO images.
@@
Accordingly, we aimed to develop and release a fully automatic analysis toolkit, SLOctolyzer, for segmentation and measurement of retinal vessels as seen on SLO images. Here, we introduce and quantitatively validate the segmentation models, trained on datasets primarily related to systemic health, to detect the en face retinal vessels into arteries and veins, and detect the optic disc and foveal pit.
@@
  • SLOctolyzer supports manual annotation to correct erroneous retinal vessel segmentation maps.
@@ 2 Methods
Fig. 1 summarises the key elements of SLOctolyzer’s pipeline, which contains a segmentation module and a measurement module (of which the latter is based on features and code produced by Automorph 13). We have released SLOctolyzer in a frictionless manner so that it can be used without author permissions, proprietary software, or specialist training. SLOctolyzer is freely available here: https://github.com/jaburke166/SLOctolyzer.
@@ 2.1 Data
Data used to build the segmentation module include RAVIR 11, a publicly available dataset of pathological SLO images with the arteries and veins segmented, one cohort of healthy volunteers and three cohorts related to systemic health: i-Test 17, a cohort of pregnant women at late gestation who are experiencing either a normative, pre-eclamptic or fetal growth restricted pregnancy; PREVENT 18; 19, a cohort of mid-life individuals, half of whom are at risk of developing later life dementia; FutureMS 20; 21, a cohort of individuals with newly diagnosed relapsing-remitting multiple sclerosis (MS). All studies and cohorts adhered to the Declaration of Helsinki and received relevant ethical approval, and informed consent from all subjects was obtained in all cases from the host institution.
@@
Table 1 summarises image characteristics of each cohort. All SLO images were captured using a spectral domain Heidelberg SPECTRALIS HRA+OCT imaging device (Heidelberg Engineering, Heidelberg, Germany) and were either macula-centred or optic disc-centred. SLO images from every study were grayscale and captured a field of view of 30-degrees at the posterior pole which approximately images a 9 mm2 region of interest, using a broadband laser light source with wavelength 820nm. All cohorts used a HRA+OCT SPECTRALIS Standard Module except for the i-Test study, which used a portable version, the SPECTRALIS FLEX, for 24 participants out of the 93 available. Given a 30-degree field of view covering approximately a 9 mm2 region of interest, the SLO images of these cohorts had either approximately 25 or 50 pixels per degree of visual angle, constituting a total imaging resolution of 768 $\times$ 768 or 1536 $\times$ 1536 (pixel height $\times$ width), which had transversal spatial sampling length-scales of approximately 11.71 microns-per-pixel or 5.85 microns-per-pixel, respectively.
@@
Study Participants Images Right eyes Retinal pathology HRA+OCT Module Image resolution [px] Location
RAVIR 11 23 23 14 Yes Standard 768 $\times$ 768 Disc
Healthy 22 7 7 7 No Standard 1536 $\times$ 1536 Disc
PREVENT 18; 19 144 285 142 No Standard 1536 $\times$ 1536 Disc
i-Test 17 93 186 93 No Standard & FLEX 768 $\times$ 768 Macula
FutureMS 20; 21 15 15 9 No Standard 1536 $\times$ 1536 Disc
@@
Table 1: Image characteristics of the five cohorts used to build SLOcolyzer’s segmentation module. Image resolution is in pixels (for both lateral and axial directions), location refers to the centring of the scan, i.e. if it’s macula-centred disc-centred. HRA, Heidelberg retina angiography; OCT, optical coherence tomography. px, pixels.
@@
SLOctolyzer’s segmentation module contains three segmentation models, one for binary vessel detection, one for fovea detection and another for artery-vein and optic disc detection. Table 2 outlines which cohorts and how many eyes/images of the cohorts each model used for training and evaluation. For all studies except for FutureMS, we selected all available participants for modelling. For FutureMS, a subset of 15 participants from the wider cohort 21 were selected. These images were chosen based on image features relevant for vessel segmentation, such as blur, illumination, contrast, and abnormal features like vessel tortuosity or optic nerve head atrophy (supplementary Fig. S2). Similarly, for the i-Test cohort, 15 participants were selected from the wider cohort for building the artery-vein-optic disc (AVOD) model.
@@ 2.2.1 Binary vessel detection
A total of 30 SLO images (RAVIR & Healthy, Table 1) were used to train and evaluate the segmentation model for binary vessel detection. 23 of these were sourced from the RAVIR dataset, of which manual pixel-level annotation of the arteries and veins were done by the same, experienced retinal image grader 11. Segmentation over the optic nerve head was only performed if the artery and vein classes were able to be resolved. Vessel labels were defined as both the artery and vein class combined. The remaining seven SLO images were made available from healthy eyes and were manually segmented by an experienced image analyst (author S.G.) using ITK-Snap 23, which supports pixel-level annotation.
@@ 2.2.2 Fovea detection
Fovea detection was performed by the same experienced grader (author J.B.) and used all 516 SLO images.
The definitions of the fovea are guided by features on the image such as brightness and shape, rather than being biologically driven. We define the foveal pit as a single pixel coordinate on the en face SLO and cross-sectional OCT B-scan at the centre of the foveola centralis. The foveola is a circular zone of approximately 350-micron width which is avascular, creating a depression at the centre of the macula which pushes the inner retinal layers laterally 24. On an OCT B-scan, the depression appears as a dip in the retina and is identified as a hyperreflective pixel at the centre of this dip, often aligning with a ridge formed at the photoreceptor layer. On the en face SLO, the depression appears as a dark spot (sometimes showing a hyperreflective dot in its centre) where smaller arterioles and venules, branching from major arteries and veins, converge to as they thin in calibre in the macula. See supplementary Fig. S3 and Fig. S4 and supporting text for exemplar en face and cross-sectional images with arrows indicating the detected foveal pit.
For the i-Test cohort, access to both the SLO and OCT was possible, which allowed the fovea coordinate on the SLO to be cross-referenced using it’s transversal position on the corresponding fovea-centred OCT B-scan. The fovea location on each OCT B-scan was detected using Choroidalyzer 16, a fully automatic toolkit for choroid analysis and fovea detection in OCT images. For the remaining images, only the SLO was available and the fovea coordinate was selected manually, using an in-house GUI written in Python (version 3.11.6).
We assessed the accuracy of single-grader fovea detection through grader J.B. re-selecting the fovea pixel coordinate for all 516 images two months after the initial manual grading, the results of which can be found in supplementary Table S1. Repeatability against the initial i-Test fovea coordinates acted as comparison to ground truth, and the remaining images were used to assess intra-rater repeatability. Intra-rater repeatability had an average intra-class correlation (ICC(3,1)) of 0.99 for both $x$ and $y$ coordinates, and comparison to ground truth scored an average ICC of 0.82 across both $x$ and $y$ coordinates.
@@ 2.2.3 Artery-vein-optic disc detection
Three experienced image analysts (authors J.B., S.G., A.T.) were each given 14 SLO images to manually annotate the arteries, veins, and optic disc. For each SLO image, a binary vessel map was generated using the binary vessel detection model. Each grader then used ITK-Snap 23 for anatomical segmentation of arteries, veins and the optic disc. See supplementary Fig. S5 and S6 for the vessel labelling rules and segmentation protocol each rater followed for consistent and comparative segmentation.
Each set of 14 images had two images with distinct overlap between pairwise graders. Thus, there were a total of 30 unique SLO images used for modelling: 15 from the i-Test cohort and 15 from the FutureMS cohort. This overlap allowed for a total of six SLO images which could be used for inter-rater assessment. We calculated inter-rater agreement for all-vessel, artery, vein, and optic disc segmentation using the Dice coefficient, whose results can be found in supplementary Table S2. All values were found to be greater than 0.93, suggesting an excellent degree of consistency between graders.
@@ 2.2.4 The RAVIR dataset
The RAVIR dataset 11 was intended to be used for training the artery-vein-optic disc (AVOD) model, as it contains suitable examples of severe retinal pathology. However, before modelling we reviewed the segmentation labels and observed inconsistent artery and vein classification, an example of which can be seen in supplementary Fig. S8. Therefore, we decided to use the RAVIR dataset as an external test set for the AVOD model, after correcting for any misclassified veins or arteries manually using ITK-Snap 23. This external test set presents a significant challenge to the AVOD model as it contains examples of severe retinal pathology.
@@ 2.3 Segmentation module
We used python 3.12, PyTorch 2.0, Segmentation Models PyTorch 25 and MONAI 26.
@@ 2.3.1 Binary vessel detection
The model and training pipeline has been discussed previously 27. Briefly, the model was a custom UNet deep learning architecture trained from scratch, with details on how to reproduce the model architecture shown in supplementary Fig. S9 with surrounding text.
The model was trained on a total of 30 SLO images. The dataset was split randomly into training (n=24), validation (n=2) and test (n=4) sets. Offline data augmentation randomly cropped each image into 20 patches of size 320 $\times$ 240 to artificially increase the training set, reduce computational load and accelerate the training rate.
The model was trained to minimise the Dice Focal loss 26, using the Adam 28 optimiser (with default optimiser hyperparameters) for 600 epochs using a batch size of 20. The initial learning rate was set as 1e-3 until epoch 300, at which point the learning rate was reduced to 1e-4 for the remainder of training.
@@ 2.3.2 Fovea detection
Fovea segmentation is a significantly easier task than vessel detection or artery-vein classification, so our model used a pre-trained 25, lightweight UNet deep learning architecture with a MobileNetV3 29 backbone, previously trained using ImageNet 30. Details on how to reproduce the model architecture are shown in supplementary code listing LABEL:suppcode:fovea_code with surrounding text.
This model used all available images from all cohorts, comprising 516 SLO images. The dataset was split into training ($n=359$), validation ($n=104$) and test sets ($n=54$) at the participant-level, so there was no overlap of participants between sets. All images were resized to a common image resolution of 768 $\times$ 768 pixels for training. To facilitate learning through segmentation, training labels were circular binary masks whose centre was the fovea coordinate with a radius of 60 pixels.
The model was trained to minimise the binary cross entropy loss using the AdamW 31 optimiser for 200 epochs using a batch size of 8. During training, we applied two sets of random augmentations per image, per epoch, artificially doubling the batch size. An initial learning rate of 5e-4 was set, and after a 10-epoch linear warm up to a learning rate of 5e-3, it decayed following a cosine relationship for 90 epochs, with the learning rate set at 5e-4 for the remaining 100 epochs.